Bart's age is one-third of his mother's age. 7 years ago, his age was one-fifth of hers. How old are both now?
Answer by mananth(16946)

Bart is 1/3 of mother's age
So mother is 3 times Bart's age
Bart's age =x
mother's age =3x
..
7 years ago
Bart is 1/5 of her age
so she is 5 times his age
bart --- x-7
mother --- 3x-7
3x-7 =5(x-7)
3x-7= 5x-35
5x-3x=35-7
2x=28
/2
x=14 Bart's age
Mother's age => 3x = 42 years